How do you solve the equation:
t-4 divided by 6= 1 1/3

Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

t-4/6=4/3
t=8/6+4/6
t=12/6
t=2

Answer by Alan3354(69443)

How do you solve the equation:
t-4 divided by 6= 1 1/3
----------------
Any way that works.
I would eliminate fractions first, it reduces the chance of error.
(t-4)/6 = 4/3
Multiply by 6
t-4 = 8
t = 12